Bernoulli’s Equation

(y^4 − 2xy)dx + 3x^2 dy = 0 when y(1)=0


1
Expert's answer
2022-03-29T17:26:19-0400

"(y^4 \u2212 2xy)dx + 3x^2 dy = 0" "y(1)=0"

"3x^2\frac{dy}{dx}-2xy=-y^4"

"u=y^{-3}"

"y'=-\frac13u^{-\frac43}u'"

"-3x^2\frac13u^{-\frac43}u'-2xu^{-\frac13}=-u^{-\frac43}"

"x^2u'+2xu=1"

"u'+2\frac {u}{x}=\frac{1}{x^2}" (1)

Let’s solve the following equation

"u'+2\frac {u}{x}=0"

"\frac {du}{u}=-2\frac{dx}{x}"

"u=C(x)x^{-2}"

To find solution of the equation (1) we should think C is a function of x

"u'=C'x^{-2}-2Cx^{-3}"

Substitution u and u’ into (1) gives

"C'x^{-2}-2Cx^{-3}+2C(x)x^{-3}=\frac{1}{x^2}"

"C'=1"

"C=x+C_1"

"y^{-3}=u=\frac{(x+C_1)}{x^2}"

"y=(\frac{x^2}{x+C_1})^{\frac13}= (\frac{C_2x^2}{C_2x+1})^{\frac13}"

"y(1)=(\frac{C_2}{C_2+1})^{\frac13}=0"

"C_2=0"

"y=0" .

Answer: "y=0"